Can you please help me on this question: 1)State whether the relation is a function and give the domain and range.
a)y=x^2+6

Answer by stanbon(75887)

State whether the relation is a function and give the domain and range.
a)y=x^2+6
-----------------
Yes, it is a function.
-----------------
Domain: All Real Numbers
Range:
Since y^2 is always >= 0, y^2+6 will always be >= 6.
So the range is "All Real Numbers >= 6.
---------------------
Cheers,
Stan H.